OpenStudy (anonymous):

7x - 6y = 19 for y ?

OpenStudy (ymendoza88):

7x-6y=19 -6y=19-7x equals \[\frac{ 19-7x }{-6 }\]

OpenStudy (ymendoza88):

for y

OpenStudy (anonymous):

You don't have to have to set it as two fractions? You just leave it as 19 - 7x / -6 ?

OpenStudy (ymendoza88):

you could set it as \[\frac{ 19 }{ -6 } - \frac{ 7x }{ -6 }\] but the first way is more simple.